# Weyl law for 1-cycles

### v1

@@ 3.6. Approximating continuous families
Now we want to extend $F^{\prime}$ to a continuous family with domain $X(q)$. For that purpose, we apply [12][Proposition 2.6] to the discrete $\varepsilon_{1}$-fine and $(N(k,p),\delta_{1})$-localized map $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$, obtaining an extension $F^{\prime}:X(q)\to\mathcal{Z}_{k}(M)$ such that
@@
From [12][Proposition 2.6] we also know that given a cell $C$ of $X(q)$ and $x,y\in C$,
@@
There exists $\eta_{0}(p,M)$ and $C(p)$ such that the following is true. Let $\eta<\eta_{0}$ and let $F:X^{p}\to\mathcal{Z}_{k}(M)$ be a continuous family without concentration of mass such that $\mathcal{F}(F(x))\leq\eta$ for every $x\in X$. Then there exists a continuous family $\tau:X^{p}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq C(p)\eta$.
In order to prove the proposition, we will need the following lemma.
@@
There exists a constant $C(p)>0$ such that the following is true. Let $F:X^{p}\to\mathcal{Z}_{k}(M)$ be a continuous $\delta$-localized family such that given a cell $C$ of $X$ with associated admissible family $\{U_{i}^{C}\}_{i\in I_{C}}$ it holds
$\sup_{x\in C}\{\operatorname{\mathbf{M}}(F(x)\llcorner(\bigcup_{i\in I_{C}}U_{i}^{C}))\}\leq\varepsilon.$
Assume in addition that $\mathcal{F}(F(x))\leq\eta$ for every $x\in X$. Then there exists a continuous family $\tau:X^{p}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=F(x)$ and
$\operatorname{\mathbf{M}}(\tau(x))\leq C(p)(\eta+\varepsilon).$
@@
We use the argument in [12][Proposition 2.6]. For $x\in X_{0}$ we define $\tau(x)$ to be any $(k+1)$-chain such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\eta$. Then we proceed skeleton by skeleton with the following inductive property. There exists a constant $C(l)$ and a continuous $\tau:X_{l}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau(x)=F(x)$ and $\operatorname{\mathbf{M}}(\tau(x))\leq C(l)(\eta+\varepsilon)$ for every $x\in X_{l}$.
@@
The inductive property holds for $l=0$ taking $C(0)=1$ by construction. Assume it holds for $l-1$. Let $C$ be an $l$-cell of $X^{p}$ and let $\{U_{i}\}_{i\in I_{C}}$ be the $\delta$-admissible family associated to $C$. Fix a vertex $v$ of $C$. We know that there exists a continuous family $z_{C}:C\to\mathcal{Z}_{k}(M)$ of cycles supported in $U_{C}=\bigcup_{i\in I_{C}}U_{i}^{C}$ such that
$F(x)=F(v)+z_{C}(x)$
and $\operatorname{\mathbf{M}}(z_{C}(x))\leq 2\varepsilon$ for every $x\in C$. Consider $\mu_{C}:C\to\mathcal{I}_{k+1}(M)$ given by
$\mu_{C}(x)=\sum_{i\in I_{C}}\operatorname{Cone}_{i}(z_{C}(x)\llcorner U_{i}^{C})$
where $\operatorname{Cone}_{i}(z)$ denotes the cone over $z$ centered at the central point of $U_{i}^{C}$. It is clear that $\partial\mu_{C}(x)=z_{C}(x)$ and $\operatorname{\mathbf{M}}(\mu_{C}(x))\leq 2\varepsilon\delta$. Observe that the map $\tau_{C}(x)=\tau(v)+\mu_{C}(x)$ verifies $\partial\tau_{C}(x)=F(x)$ and
$\operatorname{\mathbf{M}}(\tau_{C}(x))\leq C(l-1)(\eta+\varepsilon)+2\varepsilon\delta\leq(C(l-1)+2)(\eta+\varepsilon).$
Although $\partial\tau_{C}(x)=F(x)=\partial\tau(x)$ for $x\in\partial C$, $\tau_{C}|_{\partial C}$ and $\tau|_{\partial C}$ may not be the same function. But we know that the family $\Delta\tau_{C}:\partial C\to\mathcal{Z}_{k+1}(M)$ given by $\Delta\tau_{C}(x)=\tau_{C}(x)-\tau(x)$ is supported in $U_{C}$ and has mass bounded by $2(C(l-1)+1)(\eta+\varepsilon)$. Therefore we can extend it to all of $C$ by contracting radially on each $U_{i}^{C}$, keeping the same upper bound for the mass. For $x\in C$ define
$\tau(x)=\tau_{C}(x)-\Delta\tau_{C}(x).$
By definition of $\Delta\tau_{C}$, this provides an extension of $\tau$ to $C$ which verifies $\partial\tau(x)=\partial\tau_{C}(x)=F(x)$ because $\Delta\tau_{C}(x)\in\mathcal{Z}_{k+1}(M)$ for every $x$. In addition,
$\operatorname{\mathbf{M}}(\tau(x))\leq\operatorname{\mathbf{M}}(\tau_{C}(x))+\operatorname{\mathbf{M}}(\Delta\tau_{C}(x))\leq(3C(l-1)+4)(\eta+\varepsilon)$
which completes the proof of the inductive step by setting $C(l)=3C(l-1)+4$.
@@
Let $\varepsilon_{1}>0$ be a sufficiently small number to be determined later and consider a decreasing sequence $(\varepsilon_{j})_{j\in\mathbb{N}}$ with $\varepsilon_{j}\leq\frac{\varepsilon_{1}}{2^{j-1}}$. Also consider a sufficiently small decreasing sequence $(\delta_{j})_{j\in\mathbb{N}}$. We can use Theorem 3.15 to find for each $j\in\mathbb{N}$ a $\delta_{j}$-localized family $F_{j}:X^{p}(q_{j})\to\mathcal{Z}_{k}(M)$ such that $\mathcal{F}(F_{j}(x),F(x))\leq\varepsilon_{j}$ and in addition, using the no concentration of mass condition as in the proof of Theorem 3.15 have the property that given a cell $C$ of $X(q_{j})$,
$\sup_{x\in C}\{\operatorname{\mathbf{M}}(F_{j}(x)\llcorner(\bigcup_{i\in I_{C}}U_{i}^{C}))\}\leq\varepsilon_{j}.$
We also require $X(q_{j+1})$ to be a refinement of $X(q_{j})$ for each $j\in\mathbb{N}$. Notice that
$\mathcal{F}(F_{j}(x))\leq\mathcal{F}(F(x))+\mathcal{F}(F(x)-F_{j}(x))\leq\eta+\varepsilon_{j}.$
For each $j\geq 2$, let $G_{j}(x)=F_{j}(x)-F_{j-1}(x)$. Notice that
$\mathcal{F}(G_{j}(x))\leq\mathcal{F}(F_{j}(x)-F(x))+\mathcal{F}(F_{j-1}(x)-F(x))\leq\varepsilon_{j}+\varepsilon_{j-1}\leq 2\varepsilon_{j-1}$
and that $G_{j}$ is $\delta_{j-1}$-localized in $X(q_{j})$. In addition,
$\sup_{x\in C}\{\operatorname{\mathbf{M}}(G_{j}(x)\llcorner(\bigcup_{i\in I_{C}}U_{i}^{C}))\}\leq 2\varepsilon_{j-1}.$
Then by Lemma 3.49, for each $j\geq 2$ there exists a continuous family $\tau_{j}:X^{p}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau_{j}(x)=G_{j}(x)$ and $\operatorname{\mathbf{M}}(\tau_{j}(x))\leq 4C(p,M)\varepsilon_{j-1}$. By the same procedure, we can construct a family $\tau_{1}:X^{p}\to\mathcal{Z}_{k}(M)$ such that $\partial\tau_{1}(x)=F_{1}(x)$ and $\operatorname{\mathbf{M}}(\tau_{1})(x)\leq C(p,M)(\eta+2\varepsilon_{1})$. Observe that
$\sum_{j=1}^{\infty}\operatorname{\mathbf{M}}(\tau_{j}(x))$ $\leq C(p)(\eta+2\varepsilon_{1}+4\sum_{j=1}^{\infty}\varepsilon_{j})$
$\leq C(p)(\eta+2\varepsilon_{1}+4\sum_{j=1}^{\infty}\frac{\varepsilon_{1}}{2^{j-1}})$
$\leq C(p)(\eta+10\varepsilon_{1}).$
Therefore by the Weierstrass $M$-test, the series $\sum_{j=1}^{\infty}\tau_{j}(x)$ converges uniformly to a certain continuous map $\tau:X^{p}\to\mathcal{I}_{k+1}(M)$ in the flat topology. By the previous calculation, if we set $\varepsilon_{1}<\frac{\eta}{10}$ it holds that $\operatorname{\mathbf{M}}(\tau(x))\leq 2C(p)\eta$ for every $x\in X$. Notice that
$\partial\tau(x)$ $=\lim_{J\to\infty}\partial[\sum_{j=1}^{J}\tau_{j}(x)]$
$=\lim_{J\to\infty}\partial\tau_{1}(x)+\sum_{j=2}^{J}\partial\tau_{j}(x)$
$=\lim_{J\to\infty}F_{1}(x)+\sum_{j=2}^{J}F_{j}(x)-F_{j-1}(x)$
$=\lim_{J\to\infty}\partial F_{J}(x)$
@@
hence $\tau$ verifies the requirements in the proposition. ∎
@@ 4.1. Parametric isoperimetric inequality in $\mathbb{D}^{n}$
The goal of this section is to prove a parametric isoperimetric inequality for continuous families of absolute 00-cycles in the $n$-disk
@@
Denote $e=(0,0,...,0,-1)$ the south pole and $e^{\prime}=(0,0,...0,1)$ the north pole. Also denote $\mathbb{S}^{n-1}=\partial\mathbb{D}^{n}$. We will follow the strategy from [12] where such inequality was proved in the $2$-disk $\mathbb{D}^{2}$.
@@
We can think that the previous theorem provides a short filling $\tilde{G}:X\to\mathcal{I}_{1}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ of the family $\tilde{F}:X\to\mathcal{Z}_{0}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ of relative 00-cycles induced by $F$. Or alternatively, that reduces the problem of finding a short filling of $F$ to that of finding a short filling of the contractible family $F(x)-\partial G(x)$ which is supported in $\partial\mathbb{D}^{n}$; suggesting that proceeding by induction in the dimension may be a good idea to find short fillings of absolute families of 00-cycles on high dimensional manifolds. Observe that in order to actually have control over $\operatorname{\mathbf{M}}(G(x))$, we need to have a bound on the mass of $F(x)$ restricted to $\operatorname{int}(\mathbb{D}^{n})\cup B$. Notice that radius $L$ of $B$ can be arbitrarily small and we will take advantage of that in the applications of the theorem. Indeed, we will show in the next section that any $\delta$-localized family $F$ for which $\operatorname{\mathbf{M}}(F(x)\llcorner\operatorname{int}(\mathbb{D}^{n}))$ is controlled can be slightly perturbed to a new family $F^{\prime}$ for which $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner(\operatorname{int}(\mathbb{D})\cup B))$ is controlled, where the radius $L$ of $B$ can be chosen to be any number smaller than $\delta$. We state and demonstrate some lemmas which are necessary for Theorem 4.1 and then we end the section by proving it.
@@
Let $S$ be the $1$-skeleton of a unit grid in $\mathbb{R}^{n}$ and let $T$ be the dual $(n-2)$-skeleton. Let $R>0$ and let $B\subseteq S_{R}(0)=\partial B_{R}(0)$ be a closed ball. Let $\{\tilde{T}_{i}\}_{i\in I}$ be the finite collection of $(n-2)$-dimensional affine subspaces of $\mathbb{R}^{n}$ such that $\bigcup_{i\in I}\tilde{T}_{i}\cap B_{R}(0)=T\cap B_{R}(0)$ and denote $T_{i}=\tilde{T}_{i}\cap B_{R}(0)$. Then there exists a point $P\in\mathbb{R}^{n}\setminus B_{R}(0)$ and a real number $\varepsilon>0$ verifying the following:
@@
    Let $C$ be the cone tangent to $S_{R}(0)$ centered at $P$ and let $F$ be the region enclosed by $C$ and $S_{R}(0)$. Then $\partial F\cap S_{R}(0)\subseteq B$. This implies that given a line $L$ through $P$ intersecting $B_{R}(0)$, the point in $L\cap S_{R}(0)$ closest to $P$ is in $B$.
@@
    Denote $\mathscr{L}$ the set of lines through $P$ which intersect $B_{R}(0)$, which is a compact subset of $\mathbb{RP}^{n-1}$. Then for each $L\in\mathscr{L}$, $L$ intersects $N_{\varepsilon}T_{i}$ for at most (${n\choose 2}$ different values of $i$ and for such values $L\cap N_{\varepsilon}T_{i}$ is contained in a ball $B_{i}^{L}$ of radius smaller than $1$.
@@
The proof of the previous lemma is based in Larry Guth’s construction on bending planes around a skeleton which is explained in the proof of [10][Theorem 1]. In Guth’s paper, they consider families of $k$-planes perpendicular to a fixed $(n-k)$-plane $P$ in general position and here we deal with families of lines passing through a generic point $P$. All the arguments in this section extend those of Guth and Liokumovich in [12][Section 5], where they prove the Parametric Isoperimetric Inequality for 00-cycles in the $2$-disk.
@@
where two planes $T_{i}$, $T_{i^{\prime}}$ are parallel if and only if there exists $1\leq j\leq{n\choose 2}$ such that $i,i^{\prime}\in I_{j}$ (each class of planes is determined by choosing $n-2$ different elements of the canonical basis $\{e_{1},...e_{n}\}$ of $\mathbb{R}^{n}$). Given $21\leq j\leq n-$ 2 and $i,i^{\prime}\in I_{j}$ let $V_{i,i^{\prime}}^{j}$ be the $(n-1)$-dimensional plane containing $T_{i}^{j}$ and $T_{i^{\prime}}^{j}$. Denote
@@
being $V\subseteq\mathbb{R}^{n}$ a subset of Lebesgue $n$-measure 00. As the set of points $P$ satisfying (1) has nonempty interior, we can pick such $P$ which is not in $V$. This implies that given a line $L\in\mathscr{L}$ and a natural number $1\leq j\leq{n\choose 2}$, there exists at most one $i\in I_{j}$ such that $L\cap T_{i}\neq\emptyset$ and moreover for such $i$ the intersection $L\cap T_{i}$ consists of a single point $p^{L}_{i}$. Denote $\alpha_{i}^{L}$ the angle of intersection of $L$ with $T_{i}$.
@@
Observe that $\{B(L,\delta_{L}):L\in\mathscr{L}\}$ is an open cover of $\mathscr{L}$. As $\mathscr{L}$ is compact, we can pick a finite subcover $\{B(L_{k},\delta_{L_{k}}):1\leq k\leq K\}$. Now define $\varepsilon_{1}=\min\{\varepsilon_{L_{k}}:1\leq k\leq K\}$ and $\beta=\min\{\frac{\sin(\alpha_{i}^{L_{k}})}{2}:1\leq k\leq K,i\in I_{L_{k}}\}$ and set $\varepsilon=\min\{\varepsilon_{1},\frac{\beta}{2}\}$. Let $L\in\mathscr{L}$. Let $1\leq k\leq K$ be such that $L\in B(L_{k},\delta_{L_{k}})$. Then by construction
@@
We will also need [10][Lemma 2.1] which we state below. We use the notation introduced in Lemma 4.2.

### v2

@@ 3.6. Approximating continuous families
Now we want to extend $F^{\prime}$ to a continuous family with domain $X(q)$. For that purpose, we apply [GL22][Proposition 2.6] to the discrete $\varepsilon_{1}$-fine and $(N(k,p),\delta_{1})$-localized map $F^{\prime}:X(q)_{0}\to\mathcal{Z}_{k}(M)$, obtaining an extension $F^{\prime}:X(q)\to\mathcal{Z}_{k}(M)$ such that
@@
From [GL22][Proposition 2.6] we also know that given a cell $C$ of $X(q)$ and $x,y\in C$,
@@
For each $\varepsilon>0$, there exists $\eta>0$ such that the following holds. Let $F:X^{p}\to\mathcal{Z}_{k}(M)$ be a continuous family such that $\mathcal{F}(F(x))\leq\eta$ for every $x\in X$. Then there exists $\tau:X^{p}\to\mathcal{I}_{k+1}(M)$ such that $\partial\tau=F$ and $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$ for every $x\in X$.
In order to prove the proposition, we first prove the following intermediate result.
@@
For each $\varepsilon>0$, there exists $\eta>0$ such that the following holds. Let $F:X^{p}\to\mathcal{Z}_{k}(M)$ be a continuous family such that $\mathcal{F}(F(x))\leq\eta$ for every $x\in X$. Then given $\alpha>0$, there exist continuous families $F_{1},F_{2}:X^{p}\to\mathcal{Z}_{k}(M)$ and $\tau:X^{p}\to\mathcal{I}_{k+1}(M)$ such that
  1. (1)
    $F=F_{1}+F_{2}$,
  2. (2)
    $\mathcal{F}(F_{2}(x))\leq\alpha$,
  3. (3)
    $\partial\tau(x)=F_{1}(x)$,
  4. (4)
    $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$.
@@
Let $\delta>0$ be a small number. Assume without loss of generality that the cubulation of $X$ is fine enough so that $F$ is $\alpha^{\prime\prime}$-fine for $\alpha^{\prime\prime}$ sufficiently small (to be chosen in terms of $\alpha$ and $\delta$). By Theorem 3.9 and Theorem 3.10, if $\eta$ is small enough (in terms of $\varepsilon$ and $\delta$), there exist $q>0$ and monotonously $(N(k,p),\delta)$-localized families $F_{1}:X(q)_{0}\to\mathcal{Z}_{k}(M)$ and $\tau:X(q)_{0}\to\mathcal{I}_{k+1}(M)$ such that
  1. (1)
    $F_{1}(x)=F(x)$ for all $x\in X_{0}$,
  2. (2)
    $F_{1}$ is $\alpha^{\prime}$-fine in $X$,
  3. (3)
    $\partial\tau(x)=F_{1}(x)$ for all $x\in X(q)_{0}$,
  4. (4)
    $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon^{\prime}$
for $\varepsilon^{\prime}$ and $\alpha^{\prime}$ small to be determined later (in terms of $\alpha$, $\varepsilon$ and $\delta$). Given $C$ a cell of $X(q)$, denote $\{U_{i}^{C}\}_{i\in I_{C}}$ the $(N(k,p),\delta)$-admissible family in which $F_{1}|_{C}$ and $\tau|_{C}$ are localized. Now we want to extend $F_{1}$ and $\tau$ to continuous families $F_{1}:X(q)\to\mathcal{Z}_{k}(M)$ and $\tau:X(q)\to\mathcal{I}_{k+1}(M)$ respectively. We proceed skeleton by skeleton with the following inductive property.
Inductive property. There exist a constant $C(l)$ and continuous extensions $F_{1}:X(q)_{l}\to\mathcal{Z}_{k}(M)$ and $\tau:X(q)_{l}\to\mathcal{I}_{k+1}(M)$ such that
  1. (D1)
    $\partial\tau(x)=F_{1}(x)$,
  2. (D2)
    $\operatorname{\mathbf{M}}(\tau(x))\leq C(l)\varepsilon^{\prime}$ for every $x\in X(q)_{l}$,
  3. (D3)
    $F_{1}$ is $C(l)\alpha^{\prime}$-fine in $X(q)$,
  4. (D4)
    $\tau|_{E\cap X(q)_{l}}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}}$ for every cell $E$ of $X(q)$.
@@
The inductive property holds for $l=0$ taking $C(0)=1$. Assume it holds for l1l-1. Let $C$ be an $l$-cell of $X^{p}$ and fix a vertex $v$ of $C$. We know that there exist continuous families $z_{C}:\partial C\to\mathcal{Z}_{k}(M)$ and $\tau_{C}:\partial C\to\mathcal{I}_{k+1}(M)$ supported in $U_{C}=\bigcup_{i\in I_{C}}U_{i}^{C}$ such that
$F_{1}(x)$ $=F_{1}(v)+z_{C}(x),$
$\tau(x)$ $=\tau(v)+\tau_{C}(x),$
$\mathcal{F}(z_{C}(x))\leq C(l-1)\alpha^{\prime}$ and $\operatorname{\mathbf{M}}(\tau_{C}(x))\leq C(l)\varepsilon^{\prime}$ for every $x\in C$. We extend $z_{C}$ and $\tau_{C}$ to all of $C$ by contracting radially on each $U_{i}^{C}$. To be precise, we define identify $C$ with with a cone over $\partial C$ and define
$z_{C}(x,t)$ $=\sum_{i\in I_{C}}(1-t)[z_{C}(x)\llcorner U_{i}]$
$\tau_{C}(x,t)$ $=\sum_{i\in I_{C}}(1-t)[\tau_{C}(x)\llcorner U_{i}].$
Defining $F_{1}(y)=F_{1}(v)+z_{C}(y)$ and $\tau(y)=\tau(v)+\tau_{C}(y)$ for $y\in C$, we obtain extensions of $F_{1}$ and $\tau$ satisfying $\partial\tau(y)=F_{1}(y)$. In addition, if $y$ corresponds to $(x,t)$ in our previous identification then
$\operatorname{\mathbf{M}}(\tau(y))\leq\operatorname{\mathbf{M}}(\tau(v))+\operatorname{\mathbf{M}}(\tau_{C}(x,t))\leq\operatorname{\mathbf{M}}(\tau(v))+\operatorname{\mathbf{M}}(\tau_{C}(x))\leq 3C(l-1)\varepsilon^{\prime}$
and
$\mathcal{F}(F_{1}(y)-F_{1}(v))=\mathcal{F}(z_{C}(x,t))\leq\mathcal{F}(z_{C}(x))\leq C(l-1)\alpha^{\prime}.$
This implies that $F_{1}:X(q)_{l}\to\mathcal{Z}_{k}(M)$ is $3C(l-1)\alpha^{\prime}$-fine in $X(q)$ and that (D1) to (D4) follow with C(l)=3C(l1)C(l)=3C(l-1), which completes the inductive step.
Now, if we choose $\alpha^{\prime}$ such that $3C(p)\alpha^{\prime}<\alpha$, $\varepsilon^{\prime}$ such that $C(p)\varepsilon^{\prime}=\varepsilon$ and $\eta$ sufficiently small in terms of $\varepsilon$ and $\delta$, we can perform the previous construction and obtain $F_{1}$, $\tau$ and $F_{2}=F-F_{1}$ satisfying
$\mathcal{F}(F_{2}(x))\leq\mathcal{F}(F_{1}(x)-F_{1}(w))+\mathcal{F}(F_{1}(w))-F_{1}(v))+\mathcal{F}(F(v)-F(x))\leq\alpha$
and the remaining desired properties, which completes the proof.
@@
Given $\varepsilon>0$, define $\varepsilon_{i}=\frac{\varepsilon}{2^{i}}$. Let $\eta_{i}>0$ be given by Lemma 3.46 for $\varepsilon=\varepsilon_{i}$. By making them smaller if necessary, we can assume $\lim_{i\to\infty}\eta_{i}=0$. Set $\eta=\eta_{1}$ and fix $F:X^{p}\to\mathcal{Z}_{k}(M)$ with $\mathcal{F}(F(x))\leq\eta$. We inductively define functions $F_{i},G_{i}:X^{p}\to\mathcal{Z}_{k}(M)$ and $\tau_{i}:X^{p}\to\mathcal{I}_{k+1}(M)$ such that
  1. (E1)
    $G_{i-1}=F_{i}+G_{i}$ for every $i\geq 1$ (we set $G_{0}=F)$.
  2. (E2)
    $\mathcal{F}(G_{i}(x))\leq\eta_{i+1}$.
  3. (E3)
    $\partial\tau_{i}=F_{i}$.
  4. (E4)
    $\operatorname{\mathbf{M}}(\tau_{i}(x))\leq\varepsilon_{i}$.
Applying Lemma 3.46 to $F=G_{0}$, $\varepsilon=\varepsilon_{1}$, $\eta=\eta_{1}$ and $\alpha=\eta_{2}$, we can define $F_{1}$ and $G_{1}$ satisfying (E1) to (E4) above. Assuming $F_{i-1},G_{i-1}$ were defined, apply Lemma 3.46 to $F=G_{i-1}$, $\varepsilon=\varepsilon_{i}$, $\eta=\eta_{i}$ and $\alpha=\eta_{i+1}$ to obtain $F_{i}$, $G_{i}$ and $\tau_{i}$ verifying the desired properties. Now define
$\tau(x)=\sum_{i=1}^{\infty}\tau_{i}(x).$
By the Weierstrass $M$-test we can see that $\tau:X\to\mathcal{I}_{k+1}(M)$ is continuous in the flat topology and is the uniform limit of the corresponding partial sums. Therefore
$\partial\tau(x)$ $=\lim_{N\to\infty}\sum_{i=1}^{N}\partial\tau_{i}(x)$
$=\lim_{N\to\infty}\sum_{i=1}^{N}F_{i}(x)$
$=\lim_{N\to\infty}\sum_{i=1}^{N}G_{i-1}(x)-G_{i}(x)$
$=\lim_{N\to\infty}G_{0}(x)-G_{N}(x)$
$=G_{0}(x)$
@@
and in addition $\operatorname{\mathbf{M}}(\tau(x))\leq\varepsilon$ for every $x$, which completes the proof. ∎
@@ 4.1. Parametric isoperimetric inequality in $\mathbb{D}^{n}$
The goal of this section is to prove a parametric isoperimetric inequality for continuous families of absolute 0-cycles in the $n$-disk
@@
Denote $\mathbb{S}^{n-1}=\partial\mathbb{D}^{n}$, e=(0,0,,0,1)e=(0,0,...,0,-1) the south pole and $e^{\prime}=(0,0,...0,1)$ the north pole. In this section and in the rest of the paper, we consider flat cycles and chains with $\mathbb{Z}_{2}$ coefficients. We will follow the strategy from [GL22] where such inequality was proved in the $2$-disk $\mathbb{D}^{2}$.
@@
We can think that the previous theorem provides a short filling $\tilde{G}:X\to\mathcal{I}_{1}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ of the family $\tilde{F}:X\to\mathcal{Z}_{0}(\mathbb{D}^{n},\partial\mathbb{D}^{n})$ of relative 0-cycles induced by $F$. Or alternatively, that reduces the problem of finding a short filling of $F$ to that of finding a short filling of the contractible family $F(x)-\partial G(x)$ which is supported in $\partial\mathbb{D}^{n}$; suggesting that proceeding by induction in the dimension may be a good idea for the purpose of finding short fillings of absolute families of 0-cycles on high dimensional manifolds. Observe that in order to actually have control over $\operatorname{\mathbf{M}}(G(x))$, we need to have a bound on the mass of $F(x)$ restricted to $\operatorname{int}(\mathbb{D}^{n})\cup B$. Notice that radius $L$ of $B$ can be arbitrarily small and we will take advantage of that in the applications of the theorem. Indeed, we will show in the next section that any $\delta$-localized family $F$ for which $\operatorname{\mathbf{M}}(F(x)\llcorner\operatorname{int}(\mathbb{D}^{n}))$ is controlled can be slightly perturbed to a new family $F^{\prime}$ for which $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner(\operatorname{int}(\mathbb{D})\cup B))$ is controlled, where the radius $L$ of $B$ can be chosen to be any number smaller than $\delta$. We state and demonstrate some lemmas which are necessary for Theorem 4.1 and then we end the section by proving the latter.
@@
Let $S$ be the $1$-skeleton of a unit grid in $\mathbb{R}^{n}$ and let $T$ be the dual (n2)(n-2)-skeleton. Let $R>0$ and let $B\subseteq S_{R}(0)=\partial B_{R}(0)$ be a closed ball. Let $\{\tilde{T}_{i}\}_{i\in I}$ be the finite collection of (n2)(n-2)-dimensional affine subspaces of $\mathbb{R}^{n}$ such that $\bigcup_{i\in I}\tilde{T}_{i}\cap B_{R}(0)=T\cap B_{R}(0)$ and denote $T_{i}=\tilde{T}_{i}\cap B_{R}(0)$. Then there exists a point $P\in\mathbb{R}^{n}\setminus B_{R}(0)$ and a real number $\varepsilon>0$ verifying the following:
@@
    Let $C$ be the cone tangent to $S_{R}(0)$ centered at $P$ and let $\Omega$ be the region enclosed by $C$ and $S_{R}(0)$. Then $\partial\Omega\cap S_{R}(0)\subseteq B$. This implies that given a line $L$ through $P$ intersecting $B_{R}(0)$, the point in $L\cap S_{R}(0)$ closest to $P$ is in $B$.
@@
    Denote $\mathscr{L}$ the set of lines through $P$ which intersect $B_{R}(0)$, which is a compact subset of $\mathbb{RP}^{n-1}$. Then for each $\in\mathscr{L}, LL$ intersects $N_{\varepsilon}T_{i}$ for at most $n\choose 2$} different values of $i$ and for such values $L\cap N_{\varepsilon}T_{i}$ is contained in a ball $B_{i}^{L}$ of radius smaller than $1$.
@@
The proof of the previous lemma is based in Larry Guth’s construction on bending planes around a skeleton which is explained in the proof of [GuthWidthVolume][Theorem 1]. In Guth’s paper, they consider families of $k$-planes perpendicular to a fixed (nk)(n-k)-plane $P$ in general position and here we deal with families of lines passing through a generic point $P$. All the arguments in this section extend those of Guth and Liokumovich in [GL22][Section 5], where they prove the Parametric Isoperimetric Inequality for 0-cycles in the $2$-disk.
@@
where two planes $T_{i}$, $T_{i^{\prime}}$ are parallel if and only if there exists $1\leq j\leq{n\choose 2}$ such that $i,i^{\prime}\in I_{j}$ (each class of planes is determined by choosing n2n-2 different elements of the canonical basis $\{e_{1},...e_{n}\}$ of $\mathbb{R}^{n}$). Given $1\leq j\leq n-2$ and $i,i^{\prime}\in I_{j}$, let $V_{i,i^{\prime}}^{j}$ be the (n1)(n-1)-dimensional plane containing $T_{i}^{j}$ and $T_{i^{\prime}}^{j}$. Denote
@@
being $V\subseteq\mathbb{R}^{n}$ a subset of Lebesgue $n$-measure 0. As the set of points $P$ satisfying (1) has nonempty interior, we can pick such $P$ which is not in $V$. This implies that given a line $L\in\mathscr{L}$ and a natural number $1\leq j\leq{n\choose 2}$, there exists at most one $i\in I_{j}$ such that $L\cap T_{i}\neq\emptyset$ and moreover for such $i$ the intersection $L\cap T_{i}$ consists of a single point $p^{L}_{i}$. Denote $\alpha_{i}^{L}$ the angle of intersection of $L$ with $T_{i}$.
@@
Observe that $\{B(L,\delta_{L}):L\in\mathscr{L}\}$ is an open cover of $\mathscr{L}$. As $\mathscr{L}$ is compact, we can pick a finite subcover $\{B(L_{k},\delta_{L_{k}}):1\leq k\leq K\}$. Now define $\varepsilon_{1}=\min\{\varepsilon_{L_{k}}:1\leq k\leq K\}$ and $\beta=\min\{\frac{\sin(\alpha_{i}^{L_{k}})}{2}:1\leq k\leq K,i\in I_{L_{k}}\}$ and pick $\varepsilon<\min\{\varepsilon_{1},\frac{\beta}{2}\}$. Let $L\in\mathscr{L}$. Let $1\leq k\leq K$ be such that $L\in B(L_{k},\delta_{L_{k}})$. Then by construction
@@
We will also need [GuthWidthVolume][Lemma 2.1] which we state below. We use the notation introduced in Lemma 4.2.